OpenStudy (idku):

A player kicks the ball. The ball attains a velocity of v m/s at a θº with respect to ground. If its trajectory begins and ends at the ground, how far from that player does it land?                                                                                                         (ignore any drag or air resistance)

OpenStudy (idku):

I know that: \(\large \color{royalblue}{{\rm V}_{ix}={\rm V}_{i}\cos\theta~~{\rm m/s}}\) and, \(\large \color{royalblue}{{\rm a}_{x}=0~~{\rm m/s^2}}\) (The velocity in the horizontal doesn't change, because there is no acceleration, thus it turns out... ) \(\large \color{royalblue}{{\rm V}_{fx}={\rm V}_{ix}+a_{x} t}\) \(\large \color{royalblue}{\int {\rm V}_{fx}~{\rm d}t=\int{\rm V}_{ix}+a_{x} t~{\rm d}t}\) \(\large \color{royalblue}{ {\rm D}_{x}={\rm V}_{ix}t+0.5~a_{x} t^2}\) \(\large \color{royalblue}{ {\rm D}_{x}={\rm V}_{i}t\cos\theta}\) But when I given the angle of 36 degrees and initial velocity of 22.7 m/s, I am still missing the time.